# From Course to Skill: Evaluating LLM Performance in Curricular Analytics

### v1

@@ Abstract
Curricular analytics (CA) – systematic analysis of curricula data to inform program and course refinement – becomes an increasingly valuable tool to help institutions align academic offerings with evolving societal and economic demands. Large language models (LLMs) are promising for handling large-scale, unstructured curriculum data, but it remains uncertain how reliably LLMs can perform CA tasks. In this paper, we systematically evaluate four text alignment strategies based on LLMs or traditional NLP methods for skill extraction, a core task in CA. Using a stratified sample of 400 curriculum documents of different types and a human-LLM collaborative evaluation framework, we find that retrieval-augmented generation (RAG) to be the top-performing strategy across all types of curriculum documents, while zero-shot prompting performs worse than traditional NLP methods in most cases. Our findings highlight the promise of LLMs in analyzing brief and abstract curriculum documents, but also reveal that their performance can vary significantly depending on model selection and prompting strategies. This underscores the importance of carefully evaluating the performance of LLM-based strategies before large-scale deployment.
@@ 1 Background
Recent advances in large language models (LLMs) offer new possibilities for curricular analytics. Their ability to efficiently extract semantics of natural language could improve how we analyze and interpret curricular content and help identify complex educational constructs and curricular elements that were previously hard to capture. In addition, their natural language interfaces lower technical barriers, making CA more accessible to educators and researchers without extensive technical expertise. As a result, increasing efforts have been made to incorporate LLMs as analytical tools in CA research [25, 6, 16, 13, 14, 18, 17]. While these studies have shown some promise of LLM-assisted CA, the reliability and generalizability of this promise across different curricular contexts are still not well understood.
@@ 2.2 Skill Extraction
Token-Based (TF-IDF): Following [15], we calculate alignment scores between each course and DWA skill using TF-IDF weights, combined with relevance weights based on token importance in the DWA dataset relative to Wikipedia. The weighted TF-IDF scores are summed and normalized by token count to adjust for course length.
Embedding-Based (BERT): We adapt the embedding-based matching method from [10]. SBERT, a siamese network-based model known for effective sentence embeddings [20], is used to calculate alignment scores between each curriculum document and skill description via cosine similarity.
Zero-Shot Prompting (ZERO-SHOT): We use both open-source and proprietary models—GPT-4o, Llama 3.3-70B, Gemini 1.5 Pro, and Claude 3.5 Sonnet—to perform zero-shot skill extraction. Each model is prompted with the curriculum document and the predefined skill description list to perform skill extraction. The full prompt is provided in the Supplementary Information.
Retrieval-Augmented Generative (RAG): We use RAG, a strategy that improves performance by retrieving relevant external information before generation [25], as a pre-filtering step to narrow the skill pool. Specifically, we embed 2,070 skill descriptions into a vector database and retrieve the top 20 most relevant skills [^2] The number 20 aligns with the typical number of DWAs associated with each occupation in O*NET. based on cosine similarity with the curriculum document. These retrieved skills, along with the curriculum document and query, are then used to construct a structured prompt for the LLM to extract skills.

### v2

@@ Abstract
Curricular analytics (CA) – systematic analysis of curricula data to inform program and course refinement – becomes an increasingly valuable tool to help institutions align academic offerings with evolving societal and economic demands. Large language models (LLMs) are promising for handling large-scale, unstructured curriculum data, but it remains uncertain how reliably LLMs can perform CA tasks. In this paper, we systematically evaluate four text alignment strategies based on LLMs or traditional NLP methods for skill extraction, a core task in CA. Using a stratified sample of 400 curriculum documents of different types and a human-LLM collaborative evaluation framework, we find that retrieval-augmented generation (RAG) is the top-performing strategy across all types of curriculum documents, while zero-shot prompting performs worse than traditional NLP methods in most cases. Our findings highlight the promise of LLMs in analyzing brief and abstract curriculum documents, but also reveal that their performance can vary significantly depending on model selection and prompting strategies. This underscores the importance of carefully evaluating the performance of LLM-based strategies before large-scale deployment.
@@ 1 Background
Recent advances in large language models (LLMs) offer new possibilities for curricular analytics. Their ability to efficiently extract the semantics of natural language could improve how we analyze and interpret curricular content and help identify complex educational constructs and curricular elements that were previously hard to capture. In addition, their natural language interfaces lower technical barriers, making CA more accessible to educators and researchers without extensive technical expertise. As a result, increasing efforts have been made to incorporate LLMs as analytical tools in CA research [25, 6, 16, 13, 14, 18, 17]. While these studies have shown some promise of LLM-assisted CA, the reliability and generalizability of this promise across different curricular contexts are still not well understood.
@@ 2.2 Skill Extraction
2.2.1 Token-Based (TF-IDF):
Following [15], we calculate alignment scores between each course and DWA skill using TF-IDF weights, combined with relevance weights based on token importance in the DWA dataset relative to Wikipedia. The weighted TF-IDF scores are summed and normalized by token count to adjust for course length.
2.2.2 Embedding-Based (BERT):
We adapt the embedding-based matching method from [10]. SBERT, a siamese network-based model known for effective sentence embeddings [20], is used to calculate alignment scores between each curriculum document and skill description via cosine similarity.
2.2.3 Zero-Shot Prompting (ZERO-SHOT):
We use both open-source and proprietary models—GPT-4o, Llama 3.3-70B, Gemini 1.5 Pro, and Claude 3.5 Sonnet—to perform zero-shot skill extraction. Each model is prompted with the curriculum document and the predefined skill description list to perform skill extraction. The full prompt is provided in the Supplementary Information.
2.2.4 Retrieval-Augmented Generative (RAG):
We use RAG, a strategy that improves performance by retrieving relevant external information before generation [25], as a pre-filtering step to narrow the skill pool. Specifically, we embed 2,070 skill descriptions into a vector database and retrieve the top 20 most relevant skills [^2] The number 20 aligns with the typical number of DWAs associated with each occupation in O*NET. based on cosine similarity with the curriculum document. These retrieved skills, along with the curriculum document and query, are then used to construct a structured prompt for the LLM to extract skills.